# Effect Identification and Unit Categorization in the Multi-Score Regression Discontinuity Design with Application to LED Manufacturing

### v1

@@ 3.6 Inheritance
In this section, we investigate how the unit categorization changes when altering the treatment assignment $T$, e.g. replacing it with a subrule $G$ of $T$. The goal is to avoid using knowledge of potential unobserved parts of the involved decision rules when identifying the unit behavior. Particularly, we are going to relate the unit behavior of complex rules to simpler subrules. First we investigate how the categories behave under the negation operation:
@@
Indecisive items remain stable under negation of both decision rules whereas the non-change units are stable under negation of $T$ but flip when negating $D$. This is in contrast to complier/defier units that always flip. Next, we investigate the stability of the non-change categories under sub-rules. In general one has the following bounds:
@@
Let $G$ be a cutoff rule on $\mathbb{R}^{K}$ with $\operatorname{supp}(G)\subset\operatorname{supp}(T)$.
@@
To obtain similar results for the other categories, assumptions about the relation between $G$ and $T$ are required. The next statement generalizes the AND-rule and OR-rule examples (Example 3.4 and Example A.2).
@@
Let G,HG,\,H be cutoff rules on $\mathbb{R}^{K}$ with $\operatorname{supp}(G)\neq\{0\}$. Further, let $\operatorname{supp}(T)=\operatorname{supp}(G)\oplus\operatorname{supp}(H)$.
@@
      Every unit $i$ is either a complier or a alwaystaker of $G$ with respect to $T$.
In the above setting, if the sub-rule $H$ is known, we are able to identify the unit categories regarding (G,T)(G,\,T), even if $G$ is not known. Moreover, using Proposition 7 we conclude that the introduction of new score variables in cutoff-rules does not introduce any defier or indecisives. This is not too surprising since, by definition, negation is required to obtain defier (see also Proposition 5). The previous statement can be extended to the case of a general decision rule $D$ setting as follows:
@@
Thus, complier of $T\coloneqq G\,\square\,H$ with respect to $D$ are bounded by complier and non-change units of the simpler rule $G$ by Proposition 8 2c and 8 1c. This equality does not hold in general. For this set, let $D\coloneqq(I_{1}\land I_{2})\lor I_{3}$, $T\coloneqq I_{1}\land I_{3}$ and $G\coloneqq I_{1}$. Then
@@
holds. Further, Proposition 8 2a, 8 1a and 8 2b, 8 1b resemble a factorization rule. Loosely speaking, the in between part $T$ of the rule hierarchy (G,T)(G,\,T) and (T,D)(T,\,D) factors out, becoming (G,D)(G,\,D). Using Proposition 8 and 5, we can immediately draw conclusions about the defier:
@@
Combining both statements, we note that if a unit complies regarding (T,D)(T,\,D) the categorization regarding (G,T)(G,\,T) persists for (G,D)(G,\,D). Figure 3 summarizes the above findings for the AND-case. Further, according to Proposition 7, in the absence of indecisive items these transitions are exhaustive, since any unit is either a complier or nevertaker of $G$ with respect to $T$.
@@
Figure 3: All possible category changes when transitioning from the complex rule $T=G\land H$ to the simpler rule $G$ given that $\operatorname{supp}(G)\oplus\operatorname{supp}(H)=\operatorname{supp}(T)$ and that there are no indecisives. The image summarizes Propositions 6, 7, 8 and 9. The OR-case can be summarized accordingly.
@@ 4 Empirical Application
In this section, we apply our theoretical results to a real-world decision-making scenario in opto-electronic semiconductor manufacturing. In particular, we investigate the effect of rework decisions during the color-conversion process in white light emitting diode (wLED) production on the product yield. After a brief description of the application, we derive some modeling implications in accordance to the observed data. This leads to an algorithmic description of the process, which is used as a data-generating process (DGP) complementing the empirical study with simulated runs based on artificial generated data that mimics the real-world case.
@@ 4.1 Rework
The phosphor conversion is a crucial step during production of wLEDs. In order to obtain a white color, several layers of phosphor are applied to a blue light emitting ground substrate, shifting the perceived color along a conversion curve from the blue region to the white region. During this conversion step, a production lot, consisting of $784$ individual wLEDs, is processed according to the product recipe. Each wLED in the lot undergoes the same production procedure, resulting in an identical stack of phosphor layers. The overall goal is to reach the specified target color after production is finished. To achieve this, intermediate color measurements are taken after the conversion step assessing proximity to the conversion target, for selected wLEDs in the lot. The processing steps after color conversion are adjusted so that a maximum of wLEDs reach the target color, as long as they are already close to it. In order to ensure a high yield, that is a high share of wLEDs in the lot that reach the color target, a rework decision is made using these intermediate measurements. That is, if the conversion target is not reached, another correction layer of phosphor is applied to the lot. For details on the conversion process, we refer to Cho et al. (2017) and Schwarz et al. (2024).
The specific decision is made based on two scores: The distance score, $X_{D}$, measures the distance of the mean color point $C_{1}=(C_{x},\,C_{y})$ measured after the regular conversion to the conversion target $C_{T}:=(C_{T,x},\,C_{T,y})$. The yield improvement score, $X_{Y}$, is a relative measure that evaluates the quality distribution of individual chips in a lot by calculating a hypothetical scenario in which the target is ideally met by the mean of the lot. If there is high variability in within a lot, rework treatment might have an overall negative effect on the yield even if the distance score suggests a treatment. Figure 4 depicts the decision criteria in more detail.
@@
The treatment is assigned according to the cutoff rule $T=I_{D}\land I_{Y}$ in order to maximize the outcome $Y$, which is the percentage of chips found that have reached the desired target after the production has finished.
However, an examination of the empirical data reveals that the operators responsible for performing the actual treatment $D$ do not always comply with $T$ (see Figure 5). We assume that this discrepancy is due to an informational advantage regarding the $X_{Y}$ score that the final decision makers (i.e., human operators) $D$ might have and which they use to override $T$, in order to avoid possible degradation due to the additional rework layering.
We formalize this cautious operator assumption in the following section, since it aligns well with the observed one-sided fuzziness in the $I_{Y}$ dimension and the strict compliance to the distance rule $I_{D}$.
Figure 5: Real data plot w.r.t. the score components $X_{D}$ and $X_{Y}$. The decision boundary $T$ is dashed. The actual treatment assignment (red and blue) follows an unobserved rule $D$, rendering the MRD fuzzy.
A policymaker or manager might be interested in validating the decision policy $T$ in this application setting. We are using the subset identification from Section 3, to estimate effects at the cutoff. By evaluating these effects, we draw conclusions for improvements of the treatment assignment $T$.
@@ 4.2 Modeling assumptions and implications
In this section, we derive modeling assumptions on the decision rule $D$ for the application setting combining the insights from the observed data and the theory. Particularly, we assume that $D$ has an information advantage over the initial decision rule $T=I_{D}\land I_{Y}$. We make the assumption that this information advantage is due to the availability of more detailed information on the improvement score. Particularly, We assume that the initial treatment assignment $T$ is only due to the improvement estimates of every $m$-th item in the production lot resulting in the score $X_{Y}$ whereas the final decision-maker has access to the overall estimated yield improvement $X_{E}=X_{Y}+X_{R}$ including every item where $X_{R}$ is the yield improvement of items not taken into account by $X_{Y}$. See Algorithm 1 for details.
Data: seed, lot-size $n$, cutoff $c$, measurement steps $m$, yield criteria $\mathcal{Y}$, distance criteria $\mathcal{D}$, operator specific policy $D_{O}$
@@
$X_{Y}\leftarrow$ calculate improvement $\mathcal{Y}(\hat{L}_{A})-\mathcal{Y}(L)$ on every $m$-th item;
@@
$X_{E}\leftarrow$ calculate improvement $\mathcal{Y}(\hat{L}_{A})-\mathcal{Y}(L)$ on every item;
@@
The implementation of this algorithm allows us to benchmark different operator assumptions and compare it to the real-world case. In semiconductor manufacturing the term operator refers to the human decision-maker carrying out $D$. Figure 6 displays exemplary data generated due to this DGP. While the knowledgeable operator has in principle access to a better improvement estimate, there are still several ways how she might apply this information. In the following we discuss several special cases of the knowledgeable operator assumption, allowing us (in theory) to identify the unit categories.
@@
In this scenario the operator $D$ accepts whatever treatment assignment $T$ has been made, despite better knowledge. Thus, the final treatment assignment is identical to the decision rule $D=T=I_{D}\land I_{Y}$ Which amounts to the operator specific policy in Algorithm 1 to $D_{O}=I_{Y}$. Consequently, one has $\operatorname{ComP}(T,D)=\mathcal{I}$ as well as
@@
Now suppose that the operator is very considerate if it comes to an accidental degradation of the production lot due to the optional rework step. That is, whenever $T$ suggests that a rework step should be carried out, the operator may object it. This objection can be motivated due to better knowledge or further constraints (e.g. time) that the intended treatment rule $T$ does not capture. If $T$ does not suggest a rework treatment we make the assumption that the operator accepts this decision (e.g. due to the fear of possible degradation). We formalize the inclination of this negative overwrite into an additional score variable $X_{op}$ and model $D$ according to this assumption:
$D=(T\land I_{op})\lor(\overline{T}\land 0)=T\land I_{op}$
In particular we set $X_{op}=X_{E}$ and thus $D_{O}(X\,|\,c)\coloneqq I_{X_{Y}}\land I_{X_{E}}$. Thus, we have
@@
$\operatorname{ComP}(G,D)=\{i\,|\,X_{E,i}>0,\,X_{Y}>0\}\,\mbox{ and }\,\operatorname{Nt}(G,D)=\{i\,|\,X_{E,i}\leq 0\}\cup\{i\,|\,X_{Y,i}\leq 0\}$
for the subrule $G\coloneqq I_{D}$ of $T$ again using Proposition 7 or the Example 3.4. Employing Theorem 2 one can estimate the sub-set complier effect of $G$ excluding the nevertaker $\Omega\coloneqq\{i\,|\,X_{Y,i}\leq 0\}$ instead of the complier effect of $G$. Since we assume that $X_{E}$ is only known by the operator the nevertaker due to $X_{E,i}\leq 0$ not excluded. Since the condition $X_{Y,i}\leq 0$ is global, the required continuity and stability assumptions are likely to be satisfied in practical applications.
@@
Finally, imagine a operator that always uses his information advantage. That is, the reasonable operator overwrites the intended $T$ treatment using the full information $X_{E}$ instead of $X_{Y}$:
$D_{O}(X\,|\,c)=I_{E}(c)$
That means, the final decision rule $D$ can be related to $T$ by observing that:
$D(X\,|\,c)=I_{D}(c)\land\operatorname{\mathbbm{1}}[X_{Y}+X_{R}>c_{Y}]$
Thus $i\in\operatorname{ComP}(T,\,D)$ if $X_{R,i}=0$, which means that there is no overall improvement or degradation on the left out items. Now suppose that $X_{R,i}\neq 0$. Then we can choose $c_{Y}=2\max(X_{Y,i},X_{R,i})$, and $c_{D}<X_{D,i}$ to obtain $D(X_{i}\,|\,c)=T(X_{i}\,|\,c)=0$. But there exist also $c\in\operatorname{supp}(T)=\mathbb{R}^{2}$ such that $D(X_{i}\,|\,c)\neq T(X_{i}\,|\,c)$. For this set $c_{D}<X_{D,i}$. If $X_{R}\geq 0$ choose $c_{Y}=X_{Y}$ then $X_{Y}+X_{R}>c_{Y}$ and thus $T(X_{i}\,|\,c)=0$ and $D(X_{i}\,|\,c)=1$. Otherwise, choose $c=X_{Y}+X_{R}$ then $c<X_{Y}$ and thus $T(X_{i}\,|\,c)=1$ and $D(X_{i}\,|\,c)=0$. This shows that
$\operatorname{ComP}(T,\,D)=\{i\,|\,X_{R,i}=0\}$
and that every other item falls into the indecisive category. There are several shortcomings with this model. First, it does not reflect the observed data, since according to this model there should be treated items that did not get the initial treatment assignment. Second, the identification result relies on the absence of indecisive items. In the application small deviations of improvement score, that is small $X_{R}$, would also be considered compliant to the assignment $T$, requiring a local definition of the categories.
While still interesting for further theoretical investigations and as an example for the indecisive case we focus on the edge-cases “the acknowledging operator” and -“the cautious operator” as visible in Figure 6 in our investigation.
@@ 5 Simulation Study
In this section, we provide numerical results from the semi-synthetic data-generating process derived in the previous section. We aim at estimating the causal effect of the rework decision on the final yield in a neighborhood of the decision boundary. For this, we will use different estimators at both decision thresholds separately and make use of the identification theorems in Section 3.

### v2

@@ 3.6 Inheritance
In this section, we investigate how the categorization of units changes when the treatment assignment $T$ is altered, e.g., by replacing it with a subrule $G$. The goal is to identify unit behavior without relying on knowledge of unobservable parts of the decision rules. In particular, we relate unit behavior under complex rules to that under simpler subrules. We begin by considering the effect of negation:
@@
Indecisive units remain unchanged under negation of either decision rule. Non-change units are stable under negation of $T$ but flip when $D$ is negated. In contrast, compliers and defiers always flip. Next, we investigate the stability of the non-change categories under sub-rules. In general, one has the following bounds:
@@
Let $G$ be a cutoff rule on $\mathbb{R}^{K}$ with $\operatorname{supp}(G)\subset\operatorname{supp}(T)$. Then:
@@
To obtain similar results for the other categories, assumptions about the relation between $G$ and $T$ are required. The following proposition generalizes the AND-rule and OR-rule examples (Example 3.4 and Example A.2).
@@
Let G,HG,\,H be cutoff rules on $\mathbb{R}^{K}$ with $\operatorname{supp}(G)\neq\{0\}$. Suppose further that $\operatorname{supp}(T)=\operatorname{supp}(G)\oplus\operatorname{supp}(H)$.
@@
      Every unit $i$ is either a complier or an alwaystaker of $G$ with respect to $T$.
In this setting, if the subrule $H$ is known, we can identify the unit categories with respect to (G,T)(G,\,T), even if $G$ itself is not fully observed. Using Proposition 7, we conclude that introducing new score variables into cutoff rules does not create any defiers or indecisive units: by definition, negation is required for these categories (see Definition 3 and Proposition 5). The previous statement can be extended to the case of a general decision rule $D$ setting as follows:
@@
Thus, compliers of $T\coloneqq G\,\square\,H$ with respect to $D$ are bounded by the complier and non-change categories of the simpler rule $G$ as stated in Proposition 8 2c and 8 1c. However, the set equalities do not hold in general. For example, let $D\coloneqq(I_{1}\land I_{2})\lor I_{3}$, $T\coloneqq I_{1}\land I_{3}$ and $G\coloneqq I_{1}$. Then
@@
holds. Furthermore, Proposition 8 2a, 8 1a and 8 2b, 8 1b resemble a factorization rule: loosely speaking, the intermediate rule $T$ in the hierarchy (G,T)(G,\,T) and (T,D)(T,\,D) factors out, becoming (G,D)(G,\,D). Using Propositions 8 and 5, we can immediately draw conclusions about defiers:
@@
Combining both statements, we see that if a unit is a complier with respect to (T,D)(T,\,D), its categorization under (G,T)(G,\,T) also holds for (G,D)(G,\,D). Figure 3 summarizes the above findings for the AND-case. Further, according to Proposition 7, in the absence of indecisive items, these transitions are exhaustive, since any unit is either a complier or nevertaker of $G$ with respect to $T$.
@@
Figure 3: All possible category changes when transitioning from the complex rule $T=G\land H$ to the simpler subrule $G$ given that $\operatorname{supp}(G)\oplus\operatorname{supp}(H)=\operatorname{supp}(T)$ and assuming that there are no indecisive units. The figure summarizes Propositions 6, 7, 8 and 9. The OR-case can be described analogously.
@@ 4 Empirical Application
In this section, we apply our theoretical results to a real-world decision-making scenario in opto-electronic semiconductor manufacturing. We analyze the effect of rework decisions during the color-conversion process in the production of white light emitting diode (wLED) on overall product yield. After a brief description of the application, we derive modeling implications based on the observed data. This leads to an algorithmic description of the production process, which serves as a data-generating process (DGP) and complements the empirical study with simulations using artificially generated data that mimic the real-world case.
@@ 4.1 Rework
Phosphor conversion is a crucial step in wLED production. To obtain white light, several layers of phosphorous substrate are applied to a blue light-emitting semiconductor, shifting the perceived color along a conversion curve from the blue region towards the white region. In this step, each production lot, consisting of $784$ individual wLEDs, is processed according to a standardized recipe. Every wLED in the lot undergoes the same procedure, resulting in an identical stack of phosphor layers. The goal is to reach the specified target color by the end of production. To achieve this, intermediate color measurements are taken for selected wLEDs in the lot after the conversion step to assess proximity to the target. Subsequent processing steps are then adjusted to maximize the number of wLEDs that meet the target color, provided they are already close. To ensure a high yield (i.e., a large share of wLEDs in the lot reaching the target), a rework decision is made based on these intermediate measurements. If the target is not met, a correction layer of phosphor is applied to the entire lot. For further details on the conversion process, see Cho et al. (2017) and Schwarz et al. (2024).
The rework decision is based on two scores: The distance score, $X_{D}$, measures the distance between the mean color point $C_{1}=(C_{x},\,C_{y})$ after the regular conversion and the target color $C_{T}:=(C_{T,x},\,C_{T,y})$. The yield-improvement score, $X_{Y}$, is a relative measure of quality variation within a lot. It evaluates a hypothetical scenario in which the target is ideally met by the mean color of the lot. If variability within the lot is high, rework may reduce overall yield even if the distance score suggests treatment. Figure 4 illustrates the decision criteria in more detail.
@@
Treatment is assigned according to the cutoff rule $T=I_{D}\land I_{Y}$, with the goal of maximizing the outcome $Y$, defined as the percentage of chips that reach the target color by the end of production.
In practice, however, the empirical data show that the operators responsible for performing the treatment $D$ do not always comply with $T$ (see Figure 5). We attribute this discrepancy to an informational advantage regarding the $X_{Y}$ score, which human operators may use to override $T$ in order to avoid possible yield losses from applying another rework layer.
We formalize this cautious-operator assumption in the following section. It reflects the observed one-sided fuzziness in the $I_{Y}$ dimension, as well as the strict compliance with the distance rule $I_{D}$.
Figure 5: Real data plot with respect to the score components $X_{D}$ and $X_{Y}$. The decision boundary $T$ is dashed. The actual treatment assignment (red and blue) follows an unobserved rule $D$, rendering the MRD fuzzy.
From a policy or management perspective, we are interested in assessing the validity of the decision rule $T$ in this setting. To this end, we apply the subset-identification framework from Section 3 to estimate effects at the cutoff. By evaluating these effects, we draw conclusions for improving the treatment assignment $T$.
@@ 4.2 Modeling assumptions and implications
In this section, we derive modeling assumptions on the decision rule $D$ for the application setting, combining insights from the observed data and the theory. We assume that $D$ has an informational advantage over the initial decision rule $T=I_{D}\land I_{Y}$. Specifically, we posit that this advantage arises from having more detailed information about the yield-improvement score.
The initial treatment assignment $T$ is based only on the improvement estimates of every $m$-th item in the production lot, resulting in the score $X_{Y}$. In contrast, the final decision-maker has access to an overall yield-improvement estimate $X_{E}=X_{Y}+X_{R}$, where $X_{R}$ reflects the contribution of items not included in $X_{Y}$. See Algorithm 1 for details.
Data: seed, lot-size $n$, cutoff $c$, measurement steps $m$, yield criteria $\mathcal{Y}$, distance criteria $\mathcal{D}$, operator-specific policy $D_{O}$
@@
$X_{Y}\leftarrow\mathcal{Y}(\hat{L}_{A})-\mathcal{Y}(L)$ on every $m$-th item;
@@
$X_{E}\leftarrow\mathcal{Y}(\hat{L}_{A})-\mathcal{Y}(L)$ on every item;
@@
Implementing this algorithm allows us to benchmark different operator assumptions and compare them with the real-world case. In semiconductor manufacturing, the term operator refers to the human decision-maker implementing $D$. Figure 6 shows an example of data generated by this DGP.
Although the knowledgeable operator has, in principle, access to a better improvement estimate, there are several ways in which this information may be applied. In the following, we discuss special cases of the knowledgeable-operator assumption, each of which (at least in theory) permits identification of the unit categories.
@@
In this scenario, the operator $D$ accepts the treatment assignment prescribed by $T$ despite having better information. Thus, the final treatment assignment coincides with the initial decision rule: $D=T=I_{D}\land I_{Y}$. This corresponds to the operator-specific policy in Algorithm 1 with $D_{O}=I_{Y}$. Consequently, we obtain
$\operatorname{ComP}(T,D)=\mathcal{I}$
as well as
@@
Suppose the operator is particularly careful to avoid accidental degradation of the production lot caused by the optional rework step. In cases where $T$ recommends rework, the operator may override this recommendation. Such overrides may be motivated either by better information or by additional constraints (e.g., time) that the intended treatment rule $T$ does not capture. In contrast, if $T$ does not recommend rework, we assume the operator accepts this decision (e.g., out of concern about possible degradation).
We formalize this “negative overwrite” by introducing an additional score variable $X_{op}$ and modeling $D$ as:
$D=(T\land I_{op})\lor(\overline{T}\land 0)=T\land I_{op}.$
In particular, we set $X_{op}=X_{E}$, which leads to
$D_{O}(X\,|\,c)\coloneqq I_{X_{Y}}\land I_{X_{E}}.$
Thus, we have
@@
$\operatorname{ComP}(G,D)=\{i\,|\,X_{E,i}>0,\,X_{Y,i}>0\}\,\mbox{ and }\,\operatorname{Nt}(G,D)=\{i\,|\,X_{E,i}\leq 0\}\cup\{i\,|\,X_{Y,i}\leq 0\}$
for the subrule $G\coloneqq I_{D}$, again using Proposition 7 or Example 3.4.
Employing Theorem 2, one can estimate the subset complier effect of $G$, excluding the never-taker group $\Omega\coloneqq\{i\,|\,X_{Y,i}\leq 0\}$, instead of the overall complier effect of $G$. Since we assume that $X_{E}$ is known only to the operator, nevertakers defined by $X_{E,i}\leq 0$ cannot be excluded. However, because the condition $X_{Y,i}\leq 0$ applies globally, the continuity and stability assumptions required for identification are likely to hold in practice.
@@
Finally, consider an operator who always uses the additional information available. In this case, the reasonable operator overwrites the intended assignment rule $T$ and bases the final decision solely on the full information $X_{E}$ rather than on $X_{Y}$:
$D_{O}(X\,|\,c)=I_{E}(c).$
Accordingly, the final decision rule $D$ can be expressed in relation to $T$ as:
$D(X\,|\,c)=I_{D}(c)\land\operatorname{\mathbbm{1}}[X_{Y}+X_{R}>c_{Y}].$
A unit $i$ belongs to $\operatorname{ComP}(T,D)$ if and only if $X_{R,i}=0$, i.e., there is no improvement or degradation among the items excluded from $X_{Y}$. Now suppose $X_{R,i}\neq 0$. Then we can choose $c_{Y}=2\max(X_{Y,i},X_{R,i})$ and $c_{D}<X_{D,i}$ to obtain $D(X_{i}\,|\,c)=T(X_{i}\,|\,c)=0$. However, there also exist cutoff values $c\in\operatorname{supp}(T)=\mathbb{R}^{2}$ for which $D(X_{i}\,|\,c)\neq T(X_{i}\,|\,c)$. For units with $X_{R}\geq 0$, choose $c_{Y}=X_{Y}$. Then $X_{Y}+X_{R}>c_{Y}$, so $T(X_{i}\,|\,c)=0$ and $D(X_{i}\,|\,c)=1$. Otherwise, choose $c=X_{Y}+X_{R}$. Then $c<X_{Y}$, so $T(X_{i}\,|\,c)=1$ and $D(X_{i}\,|\,c)=0$.
This shows that
$\operatorname{ComP}(T,D)=\{i\,|\,X_{R,i}=0\},$
and that all other units fall into the indecisive category.
This model has several shortcomings. First, it does not reflect the observed data: according to the model, some treated items should not have received the initial assignment. Second, the identification result depends on the absence of indecisive items. In practice, small deviations in the improvement score (small $X_{R}$) would likely be considered compliant with the assignment $T$, which suggests the need for a more local definition of categories.
While this case is theoretically interesting (and useful as an example of the indecisive case), our empirical investigation focuses on the two edge cases: the acknowledging operator and the cautious operator, as shown in Figure 6.
@@ 5 Simulation Study
In this section, we present numerical results based on the semi-synthetic data-generating process (DGP) derived in the previous section. The aim is to estimate the causal effect of rework decisions on the final yield in a neighborhood of the decision boundary. To this end, we use different estimators at both decision thresholds separately and draw on the identification theorems from Section 3.